OpenStudy (richyw):

Let \((u,v)=\mathbf{f}(x,y,x)=(2x+y^3,xe^{5y-7z})\) Compute \(D\mathbf{f}(x,y,z),\text{ and }\partial (u,v)/\partial(x,y)\)

OpenStudy (richyw):

so I have \[D\mathbf{f}(x,y,z)=\left[\begin{matrix} \partial u / \partial x & \partial u / \partial y & \partial u / \partial z \\ \partial v / \partial x & \partial v / \partial y & \partial v/ \partial z \end{matrix}\right]\]

OpenStudy (richyw):

but I am struggling with how to find \[\frac{\partial(u,v)}{\partial(x,y)}=(10x-3y^2)e^{5y-7z}\]

OpenStudy (richyw):

@amistre64 if you could take a look at this I would be really stoked!

OpenStudy (anonymous):

are we are talking about vector differentiation here?

OpenStudy (richyw):

yes

OpenStudy (richyw):

ok so \[D\mathbf{F}(x,y,z)=\left[\begin{matrix} 2 & 3y^2 & 0 \\ e^{5y-7z} & 5xe^{5y-7z} & -7xe^{5y-7z}\end{matrix}\right]\]

OpenStudy (richyw):

so it's quite clear that the answer is the determinant if I ignored the third column.

OpenStudy (anonymous):

If the understanding of notations is correct, that is your \(\partial(u,v)\over\partial(x,y)\). D will be its Jacobian, a vector

OpenStudy (anonymous):

did I understand it correctly?

OpenStudy (richyw):

\(\partial(u,v)\over\partial(x,y)\) is scalar

OpenStudy (anonymous):

\[D=\left({\partial\over\partial x}\hat{i}+{\partial\over\partial y}\hat{j}+{\partial\over\partial z}\hat{k}\right)\]

OpenStudy (anonymous):

if you take dot product of D with (u,v), you get a scalar if you take cross product of D with (u,v), you get a vector

OpenStudy (richyw):

wait, why is the jacobian a vector?

OpenStudy (richyw):

isn't it the matrix I wrote down?

OpenStudy (anonymous):

yes.. now it needs the i, j, k row on the top to classify it as a vector.

OpenStudy (richyw):

\[D=\left[\begin{matrix} i & j & k \\2 & 3y^2 & 0 \\ e^{5y-7z} & 5xe^{5y-7z} & -7xe^{5y-7z}\end{matrix}\right]\]

OpenStudy (anonymous):

now, make it a determinant and evaluate you'd get an (i,j,k) vector

OpenStudy (richyw):

which represents what?

OpenStudy (anonymous):

rate of change of "f" along each axis

OpenStudy (anonymous):

one is "gradient" -> scalar one is "curl" -> vector

OpenStudy (richyw):

\[\det(D)=\left\langle 10\,x\,{e}^{5\,y-7\,z}-3\,{y}^{2}\,{e}^{5\,y-7\,z},-21\,x\,{y}^{2}\,{e}^{5\,y-7\,z}, -14\,x\,{e}^{5\,y-7\,z}\right\rangle\]

OpenStudy (anonymous):

yes

OpenStudy (richyw):

so my answer is just the one component of that vector.

OpenStudy (anonymous):

the entire vector would be the answer

OpenStudy (richyw):

it's not though.

OpenStudy (anonymous):

you i,j,k components in the vector notation are not in proper order

OpenStudy (anonymous):

\[\left<-21xy^2e^{5y-7z},-14x.....\right>\]

OpenStudy (richyw):

oh yeah I see that mistake I just threw it in from maxima, still, my book tells me that \[\frac{\partial (u,v)}{\partial (x,y)} = (10-3y^2)e^{5y-7z}\]\[\frac{\partial (u,v)}{\partial (y,z)} = -21xy^2e^{5y-7z}\]\[\frac{\partial (u,v)}{\partial (x,z)} = -14xe^{5y-7z}\]

OpenStudy (anonymous):

exactly.. k, i and j components respectively

OpenStudy (richyw):

so I can see that I take the term that is not in the partial derivative is the answer. but honestly I have no idea what I am doing. very frustrating.

OpenStudy (anonymous):

you mean the maxima interpreter?

OpenStudy (anonymous):

I am lost with the symbols you are using. If you can provide me with their meaning/ book sentence .. I can help

OpenStudy (richyw):

no sorry I meant I just used maxima to compute the determinant and I copied it down in the wrong order.

OpenStudy (richyw):

and btw I really appreciate your help electrokid. I just find that less people look at questions when you "medal" them and I still would like others to look at this. I wish we could medal more than one person...